OpenStudy (anonymous):

(36x a - 30x b) ÷ (-6x 2) what is the quotient

hero (hero):

Write as \[\frac{36xa - 30xb}{-6x^2}\] Then split into two fractions: \[-\frac{36ax}{6x^2} + \frac{30xb}{6x^2}\] Simplify from there

hero (hero):

Or you can take this approach: Factor the numerator first: \(−\dfrac{6x(6a−5b)}{6x^2}\) Then use cancelling techniques to simplify